# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ III.1 Case-I: Rashba nanowire with only linear SOC
The behavior of the supercurrent density $J(q)$ and the efficiency ($\eta$) of the SDE in the case when only linear Rashba SOC is present for different system parameters are shown in Fig. 4. As shown in Fig. 4(a), the supercurrent density is symmetric with respect to $q$ for $B_{z}=0$ even with $B_{x}\neq 0$ Legg et al. (2022). We note that as we start increasing the magnitude of $B_{z}$ with $B_{x}\neq 0$, the $J(q)$ becomes more and more asymmetric, leading to nonreciprocal critical supercurrents, $|J_{c}^{+}|\neq|J_{c}^{-}|$ and hence to finite SDE. This is a direct consequence of the fact that the presence of FFLO state and the asymmetry in critical momenta: $|q_{c}^{+}|\neq|q_{c}^{-}|$ is distinctly realized only in presence of both the fields $B_{x},B_{z}$ as noted in Fig. 2(a).
@@
The diode efficiency ($\eta$) as a function of $B_{x}$ for fixed values of $B_{z}$ and as a function of $B_{z}$ for fixed values of $B_{x}$ are shown in Fig. 4(b) and Fig. 4(c) respectively. In both cases, $\eta$ increases linearly at low field strengths, reaches a maximum at moderate values, and then decreases to zero as the fields intensify further. This behavior can be understood as a consequence of a competition between the asymmetry in the band structure of the system and changes in the superconducting pairing susceptibility as the magnetic field strengths are varied, consistent with previous studies Yuan and Fu (2022). Note that, the self-consistent BCS gap vanishes for very large values of the Zeeman fields $B_{x}$ and $B_{z}$ consequently leading to zero efficiency, however the resultant nanowire spectrum might still be gapped like an ordinary insulator.
@@
Preceding investigations on a Rashba nanowire featuring only linear Rashba SOC Legg et al. (2022) reported maximum diode efficiency of $\sim 2\%$, without taking into consideration self-consistency in the superconducting order parameter and fixing it to a particular value. In our study, we have performed an extensive analysis of the full parameter space self-consistently and demonstrated that the diode efficiency in the true superconducting ground state of the device can be increased to $\eta\gtrsim 45\%$ [see Fig. 4(b, c)]. The maximum values of the Zeeman fields ($B_{x},B_{z}$) were restricted to the point where the self-consistent BCS gap vanishes. We note that in this case $B_{y}$ also plays a similar role as $B_{z}$ due to the geometry of the setup (Fig. 1(a)) under consideration and hence further analysis considering $B_{y}\neq 0$ is not presented. However, $B_{y}$ plays a significant and distinct role when we consider the effect of cubic Rashba SOC terms as discussed in the next subsection.
Figure 6: Diode efficiency phase diagram with both linear and higher order SOC: The variation of the diode efficiency $\eta$ is shown in the plane of cubic Rashba SOC parameters $\delta-\gamma$ choosing “optimal” values of the Zeeman fields $(B_{x}/\Delta_{0},B_{z}/\Delta_{0})=(0.4,0.0)$ in panel (a) and $(B_{x}/\Delta_{0},B_{z}/\Delta_{0})=(0.4,0.5)$ in panel (b), respectively. We note a small enhancement in the maximum value of $\eta$ for finite $B_{z}$. The other system parameters are chosen as: $(B_{y},\alpha,\mu/\Delta_{0},\beta^{-1})$=(0(0( 0, $1$ eV-Å, $0.5$, $0.1$ meV).
@@ III.2 Case-II: Rashba nanowire with both linear and higher order SOC
The variation of the supercurrent density $J(q)$ and the diode efficiency ($\eta$) in the case when both the linear ($\alpha\neq 0$) and cubic Rashba SOC terms ($\delta,\gamma\neq 0$) are present for different system parameters are shown in Fig. 5. From the variation of $J(q)$ shown in Fig. 5(a), we note that in this case even with $B_{z}=0$ as long as $B_{x}\neq 0$ the supercurrent becomes asymmetric leading to nonreciprocal critical supercurrents, $|J_{c}^{+}|\neq|J_{c}^{-}|$ and hence to finite SDE. The variation of $\eta$ with the magnetic fields are shown in Fig. 5(b, c). We note from Fig. 5(b) that the diode efficiency $\eta$ at a fixed $B_{z}$ exhibits a sharp linear increase as a function of $B_{x}$ for small values of $B_{x}\ll\Delta$, eventually attaining a maxima around $B_{x}/\Delta_{0}=0.4$ similar to the only linear SOC case described earlier. The maximum efficiency realized increases as one increases the magnitude of $B_{z}$. However, $\eta$ shows only marginal enhancement at a fixed value of $B_{x}$ when $B_{z}$ is varied as depicted in Fig. 5(c).
@@
Figure 7: (a) The diode efficiency $\eta$ is shown as a function of $B_{y}$ (with fixed $B_{x}/\Delta_{0}=0.4$) for our system in consideration with cubic Rashba SOC: $(\delta,\gamma)=100$ eV-Å3. (b) The $\eta$ as a function of $B_{x}$ is shown with linear SOC and with only one cubic Rashba SOC term $\delta=100$ eV-Å3 but $\gamma=0$ for different values of $B_{z}$ taking $B_{y}=0$. We have chosen the model parameters as: $(\alpha,\mu/\Delta_{0},\beta^{-1})$=($1$ eV-Å, $0.5$, $0.1$ meV).
@@
Modifying the values of $\delta$ and $\gamma$ evidently changes the band dispersions of the system and consequently the efficiency of the superconducting diode will also change. We depict the diode efficiency phase diagrams in the $\delta-\gamma$ plane in Fig. 6. We obtain the phase diagrams by varying $U$ for every configuration of ($\delta,\gamma$) such that the self-consistent BCS gap $\Delta_{0}$ remains fixed. For the $B_{z}=0$ and $B_{x}\neq 0$ case, we can achieve a peak diode efficiency of approximately $13.6\%$ for an optimal choice of higher-order Rashba SOC parameters as shown in Fig. 6(a). Furthermore, the application of an additional $B_{z}$ field readily elevates $\eta$ to approximately $19.6\%$, as illustrated in Fig. 6(b).
Note that the results discussed so far were obtained by keeping $B_{y}=0$. Now, we would like to consider the effect of finite $B_{y}$ as well. In this case, $B_{y}$ serves the role of both $B_{x}$ and $B_{z}$ in realizing a finite SDE in our system by creating both gap and asymmetry in the band structure. We note that the application of $B_{y}$ produces linear increase in $\eta$ with increasing $B_{y}$ for fixed $B_{x}$ and $B_{z}$ as shown in the Fig. 7(a).
Finally, as a special case, we consider the case when $\gamma=0$ but $\alpha,\delta\neq 0$. We can still obtain a finite diode efficiency with a maximum of $12\%$ purely in presence of a finite $B_{x}$ even when $B_{z}=0$ [see Fig. 7(b)]. The inclusion of a single cubic Rashba SOC parameter $\delta$, in a system featuring linear Rashba SOC ($\alpha$), is sufficient to manifest SDE with exclusively finite $B_{x}$ (in contrast to the case where the system features only linear Rashba SOC, both $B_{x}$ and $B_{z}$ are equally necessary in realizing SDE). By further application of $B_{z}$ the diode efficiency $\eta$ is readily enhanced upto $\sim 21\%$ as demonstrated in Fig. 7(b).
@@
We emphasize that the presence of highly asymmetric behavior in the supercurrent, given a specific set of system parameters, does not necessarily guarantee a corresponding high level of non-reciprocity in the super-currents. As evident from Fig. 6(a), the system might exhibit nearly negligible diode effect ($\eta\sim 0$) even though the parameters are set such that it supports asymmetry in supercurrent.
@@ IV Summary and conclusions
This article analyzes how the interplay between linear and cubic Rashba SOC terms, combined with external Zeeman fields, can optimize the efficiency of SDE in a proximitized one-dimensional nanowire device. Using self-consistent mean-field analysis, we demonstrate that SDE, characterized by nonreciprocal critical currents ($|J_{c}^{+}|\neq|J_{c}^{-}|$), arises from Cooper pair momentum asymmetry in the helical superconducting ground state of this time-reversal and inversion symmetry broken system. Remarkably, we achieve a diode efficiency $\eta\gtrsim 45\%$ with only linear Rashba SOC significantly higher than the previously reported efficiency of $\eta\sim 2\%$ from non-self-consistent analyses Legg et al. (2022). In this case with only linear SOC, at least two magnetic fields are required to obtain a finite diode efficiency: one along the SOC direction and one perpendicular to it. This requirement stems from the fact that linear SOC effects can be eliminated through a spin-dependent gauge transformation Braunecker et al. (2010), resulting in a conventional BCS superconducting state with zero Cooper pair momentum when only one Zeeman field is present. However, the presence of two orthogonal Zeeman fields creates a spiral magnetic field texture via the gauge transformation, leading to momentum-dependent band asymmetry and asymmetric FFLO Cooper pairs Braunecker et al. (2010). In contrast, higher-order SOC effects cannot be eliminated via gauge transformations, allowing for finite diode efficiency with only one Zeeman field. However, including higher-order SOC doesn’t improve the maximum diode efficiency (only $\eta\sim 20\%$), this value remains significant compared to previously reported diode efficiencies Legg et al. (2022); Nadeem et al. (2023). Our results underscore the importance of self-consistently determining the order parameter of the superconducting ground state to fully understand the capability and optimize Rashba nanowire devices as superconducting diodes.

### v2

@@ III.1 Case-I: Rashba nanowire with only linear SOC
In the presence of only linear Rashba SOC, we uncover the SDE by investigating the behavior of the supercurrent density ($J(q)$) and the diode efficiency ($\eta$) for different system parameters as presented in Fig. 4. Fig. 4(a) shows that the supercurrent density is symmetric with respect to $q$ for $B_{z}=0$ even with $B_{x}\neq 0$ Legg et al. (2022). We note that as we start increasing the magnitude of $B_{z}$ with $B_{x}\neq 0$, $J(q)$ becomes more and more asymmetric, leading to nonreciprocal critical supercurrents, $|J_{c}^{+}|\neq|J_{c}^{-}|$ and hence to finite SDE. This is a direct consequence of the fact that the presence of FFLO state and the asymmetry in critical momenta: $|q_{c}^{+}|\neq|q_{c}^{-}|$ is distinctly realized only in presence of both the fields $B_{x},B_{z}$ as noted in Fig. 2(a).
@@
The diode efficiency as a function of $B_{x}$ for fixed values of $B_{z}$ and as a function of $B_{z}$ for fixed values of $B_{x}$ are shown in Fig. 4(b) and Fig. 4(c) respectively. In both cases, $\eta$ increases linearly at low field strengths, reaches a maximum at moderate values, and then decreases to zero as the field strengths continue to intensify. This behavior can be understood as a consequence of a competition between the asymmetry in the band structure of the system and changes in the superconducting pairing susceptibility as the magnetic field strengths are varied, consistent with previous studies Yuan and Fu (2022); Turini et al. (2022). Note that, the self-consistent BCS gap vanishes for very large values of the Zeeman fields $B_{x}$ and $B_{z}$ consequently leading to zero efficiency, however the resultant nanowire spectrum might still be gapped like an ordinary insulator.
@@
Preceding investigations on a Rashba nanowire featuring only linear Rashba SOC Legg et al. (2022) reported maximum diode efficiency of $\sim 2\%$, where the superconducting order parameter has been fixed to a particular value. Contrastingly, our study reveals that the self-consistently derived superconducting order parameter within the full parameters space has significant influence on the performance of the diode where the highest $\eta$ value is found to be $\gtrsim 45\%$. This is evident from calculation presented in Fig. 4(b,c). The vanishing of the superconducting gap within our self-consistent treatment necessitates restricting the maximum values of the magnetic fields used in our model. We note that in this case $B_{y}$ also plays a similar role as $B_{z}$ due to the geometry of the setup (see Fig. 1(a)) under consideration and hence further analysis considering $B_{y}\neq 0$ is not presented. However, $B_{y}$ plays a significant and distinct role when we consider the effect of cubic Rashba SOC terms as discussed in the next subsection.
@@ III.2 Case-II: Rashba nanowire with both linear and higher order SOC
The variation of $J(q)$ and $\eta$ in the case when both the linear ($\alpha\neq 0$) and cubic Rashba SOC terms ($\delta,\gamma\neq 0$) are present in the nanowire are shown in Fig. 5. First of all, comparing Fig. 5(a) with Fig. 4(a), we note that the cubic Rashba SOC terms interestingly lead to asymmetry in $J(q)$ even with $B_{z}=0$ as long as $B_{x}\neq 0$. This leads to nonreciprocal critical supercurrents, $|J_{c}^{+}|\neq|J_{c}^{-}|$ with a single magnetic field $B_{x}\neq 0$ in contrast to only linear SOC case discussed earlier. The variation of $\eta$ with the magnetic fields $B_{x}$ and $B_{z}$ with $B_{y}=0$ are shown in Fig. 5(b, c). We note from Fig. 5(b) that the diode efficiency at a fixed $B_{z}$ exhibits a linear increase as a function of $B_{x}$ for small values of $B_{x}\ll\Delta_{0}$, eventually attaining a maxima around $B_{x}/\Delta_{0}=0.4$ similar to the case with only linear Rashba SOC described earlier. This behavior, as in the earlier linear SOC case Yuan and Fu (2022); Turini et al. (2022), results from two competing effects: the system’s asymmetric band structure versus changes in superconducting pairing susceptibility as magnetic field strength varies. The maximum efficiency increases as one increases the magnitude of $B_{z}$. However, $\eta$ shows only marginal enhancement at a fixed value of $B_{x}$ when $B_{z}$ is varied as depicted in Fig. 5(c).
Figure 6: Diode efficiency phase diagram with both linear and higher order SOC: The variation of the diode efficiency $\eta$ is shown in the plane of cubic Rashba SOC parameters $\delta-\gamma$ choosing “optimal” values of the Zeeman fields $(B_{x}/\Delta_{0},B_{z}/\Delta_{0})=(0.4,0.0)$ in panel (a) and $(B_{x}/\Delta_{0},B_{z}/\Delta_{0})=(0.4,0.5)$ in panel (b), respectively. We note a small enhancement in the maximum value of $\eta$ for finite $B_{z}$. The other system parameters are chosen as: $(B_{y},\alpha,\mu/\Delta_{0},\beta^{-1})$=(0(0( 0, $100$ meV-nm, $0.5$, $0.1$ meV).
@@
Figure 7: The diode efficiency $\eta$ as a function of $B_{y}$ with both linear and cubic Rashba SOC terms for $B_{x}=0$ and finite $B_{x}/\Delta_{0}=0.3$ are shown in (a) and (b) respectively. The chosen model parameters are $(\alpha,\delta,\gamma,\mu/\Delta_{0},\beta^{-1})$=($100$ meV-nm, $10$ eV-nm3, $10$ eV-nm3, $0.5$, $0.1$ meV).
@@
Modifying the values of $\delta$ and $\gamma$ evidently changes the band dispersions in the normal and superconducting state of the system and consequently the efficiency of the superconducting diode will also change. We depict the diode efficiency phase diagrams in the $\delta-\gamma$ plane in Fig. 6. We obtain the phase diagrams by varying $U$ for every configuration of ($\delta,\gamma$) such that the self-consistent BCS gap $\Delta_{0}$ remains fixed. For the $B_{z}=0$ and $B_{x}\neq 0$ case, we can achieve a peak diode efficiency of approximately $15\%$ for an optimal choice of higher-order Rashba SOC parameters as shown in Fig. 6(a). We emphasize that this behavior is qualitatively different from the linear SOC case, where the efficiency is zero with only the $B_{x}$ field. This distinctive response, arising solely from higher-order SOC terms, provides a method to detect and quantify the strength of these terms in a nanowire device.
@@
The application of an additional magnetic field $B_{z}$ readily elevates $\eta$ to approximately $20\%$, as illustrated in Fig. 6(b) since effects of both the linear and cubic SOC terms are relevant. Note that the presence of highly asymmetric behavior in the supercurrent, given a specific set of system parameters, does not necessarily guarantee a corresponding high level of non-reciprocity in the super-currents. As is evident from Fig. 6(a,b), the system might exhibit nearly negligible diode effect ($\eta\sim 0$) even though the parameters are set such that there is asymmetry in supercurrent (top right corners of Fig. 6(a,b)).
Note that the results discussed so far were obtained by fixing $B_{y}=0$. The effect of finite $B_{y}$ in $\eta$ is shown in the Fig. 7. In this case, $B_{y}$ serves the role of both $B_{x}$ and $B_{z}$ in realizing a finite SDE by creating both gap and asymmetry in the band structure. Fig. 7(a) and Fig. 7(b) show that the application of $B_{y}$ produces systematic increase in $\eta$ with increasing $B_{y}$ for different values of $B_{z}$ with $B_{x}=0$ and $B_{x}\neq 0$ respectively. However, $B_{y}$ can not increase the efficiency beyond a certain limit because the superconducting state becomes unstable after some value of $B_{y}$ when the other magnetic fields are present as well.
Figure 8: The diode efficiency ($\eta$) with linear SOC and with cubic SOC of different type where $\delta=10$ eV-nm3 but $\gamma=0$ as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) are shown in (a) and (b) respectively. Other parameters are: $(\alpha,\mu/\Delta_{0},B_{y},\beta^{-1})$=($100$ meV-nm, $0.5$, 00, $0.1$ meV).
For completeness, we consider a special case with a different type of the higher order Rashba SOC terms hso when only $\delta\neq 0$ and $\gamma=0$ (with $\alpha\neq 0$ as usual). The variation of the corresponding diode efficiencies with the magnetic fields are shown in Fig. 8. A finite diode efficiency with a maximum of $\sim 12\%$ even in presence of a finite $B_{x}$ only is obtained as shown in Fig. 8(a) which is also in contrast to the linear SOC case. Thus the presence of even a single cubic Rashba parameter $\delta$ is sufficient to manifest SDE. By further application of $B_{z}$ the diode efficiency $\eta$ is readily enhanced upto $\sim 21\%$ as a function of $B_{x}$ shown in Fig. 8(a). There is only a marginal increase in efficiency by increasing $B_{z}$ for fixed $B_{x}$ as shown in the Fig. 8(b).
@@ IV Summary and conclusions
This article analyzes how the interplay between linear and cubic Rashba SOC terms, combined with external Zeeman fields, can optimize the efficiency of SDE in a proximitized one-dimensional nanowire device. Using self-consistent mean-field analysis, we demonstrate that SDE, characterized by nonreciprocal critical currents ($|J_{c}^{+}|\neq|J_{c}^{-}|$), arises from Cooper pair momentum asymmetry in the helical superconducting ground state of this time-reversal and inversion symmetry broken system. Remarkably, we achieve a diode efficiency $\eta\gtrsim 45\%$ with only linear Rashba SOC significantly higher than the previously reported efficiency of $\eta\sim 2\%$ from non-self-consistent analyses Legg et al. (2022). In this case with only linear SOC, at least two magnetic fields are required to obtain a finite diode efficiency: one along the SOC direction and one perpendicular to it. This requirement stems from the fact that linear SOC effects can be eliminated through a spin-dependent gauge transformation Braunecker et al. (2010), resulting in a conventional BCS superconducting state with zero Cooper pair momentum when only one Zeeman field is present. However, the presence of two orthogonal Zeeman fields creates a spiral magnetic field texture via the gauge transformation, leading to momentum-dependent band asymmetry and asymmetric FFLO Cooper pairs Braunecker et al. (2010). In contrast, higher-order SOC effects cannot be eliminated via gauge transformations, allowing for finite diode efficiency even with only one Zeeman field unlike the linear SOC case and this behavior can be used to detect the presence of higher-order SOC in nanowires. However, including higher-order SOC doesn’t improve the maximum diode efficiency (only $\eta\sim 20\%$), this value remains significant compared to previously reported diode efficiencies Legg et al. (2022); Nadeem et al. (2023). Our results underscore the importance of self-consistently determining the order parameter of the superconducting ground state to fully understand the capability and optimize Rashba nanowire devices as superconducting diodes.